In right triangle ABC the legs are 6 inches and 8 inches find the length of the hypotenuse.
Answer by bucky(2189)

This is a Pythagorean theorem problem. The Pythagorean theorem says that in a right triangle
the sum of the squares of the two legs equals the square of the hypotenuse. In equation form
this can be written as:
.
A^2 + B^2 = H^2
.
where A is one of the two legs, B is the other leg, and H is the hypotenuse.
.
For this problem you can set A = 6 inches and B= 8 inches. If you substitute those values into
the Pythagorean theorem equation you get:
.
6^2 + 8^2 = H^2
.
But 6^2 equals 36 and 8^2 equals 64. Therefore the equation becomes:
.
36 + 64 = H^2
.
Adding the two values on the left side reduces the equation to:
.
100 = H^2
.
You can now solve for H by taking the square root of both sides. On the left side the
square root of 100 equals 10 and on the right side the square root of H^2 is just H. So
the answer becomes 10 inches = H.
.
The hypotenuse is 10 inches long.
.
Hope this helps you to understand the problem and to see a procedure for working it through
to get an answer.
.